If a ball that is 4 meters above the ground is thrown horizontally at 4 m/s, how long will it take the ball to hit the ground?


1
Expert's answer
2021-05-17T13:32:05-0400

The time it will take to reach the ground is


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2·4}{9.8}}=0.9\text{ s}."

The time does not depend on the horizontal speed.